Let F(x,y)=2x^2+3sin y and let G(x,y)=e^x-2y.Compute the Jacobian determinant d(F ,G)/d(x,y)

1
Expert's answer
2022-02-14T16:39:48-0500

Solution:

"F(x,y)=2x^2+3\sin y 
\\G(x,y)=e^x-2y"

"\dfrac{\partial (F,G)}{\partial (x,y)}=\begin{vmatrix} \dfrac{\partial F}{\partial x}&\dfrac{\partial F}{\partial y}
\\ \dfrac{\partial G}{\partial x}&\dfrac{\partial G}{\partial y}\end{vmatrix}"

"=\begin{vmatrix} \dfrac{\partial (2x^2+3\sin y)}{\partial x}&\dfrac{\partial (2x^2+3\sin y)}{\partial y}
\\ \dfrac{\partial (e^x-2y)}{\partial x}&\dfrac{\partial (e^x-2y)}{\partial y}\end{vmatrix}"

"=\begin{vmatrix} 4x&3\cos y
\\ e^x&-2\end{vmatrix}
\\=-8x-3e^x\cos y"